form the partial equation by eliminating the function f from z=f(y/x)

1
Expert's answer
2022-04-17T08:50:09-0400

The equations can be obtained in different ways. For example, we can consider derivatives: "z_y=f'\frac{1}{x}" and "z_x=-\frac{y}{x^2}f'". Then, the following equation can be made: "z_y+\frac{x}{y}z_x=0." It is equivalent to: "yz_y+xz_x=0", except the point "y=0."